If cos(πœƒ ) = 3/5 and tan( πœƒ ) < 0 , find the exact value of each of the remaining trigonometric functions of πœƒ


1
Expert's answer
2022-05-24T17:41:43-0400

"\tan\theta=\frac{\sin\theta}{\cos\theta}<0" and "\cos\theta=\frac 35>0" so we can conclude "\sin\theta<0"

From "\sin^2\theta+\cos^2\theta=1" we can write

"\sin\theta=-\sqrt{1-\cos^2\theta}=-\sqrt{1-9/25}=-4/5"

"\tan\theta=\frac{\sin\theta}{\cos\theta}=\frac35:(-\frac45)=-\frac34"

"\cot\theta=\frac{1}{\tan\theta}=-\frac43"

"\sec\theta=\frac{1}{\cos\theta}=\frac53"

"\csc\theta=\frac{1}{\sin\theta}=-\frac54"